OpenStudy (unklerhaukus):

\[{dy \over dx}={4y^2 \over x^2}+5{y \over x}+1\]

OpenStudy (turingtest):

bookmark*

OpenStudy (unklerhaukus):

i think it is a Homogeneous Equation requiring a change of variables

OpenStudy (unklerhaukus):

(a linear shift

OpenStudy (turingtest):

I saw your work on it yesterday and saw no mistakes, so I'm curious as well

OpenStudy (akshay_budhkar):

i believe we have to solve using y=mx where m= y/x

OpenStudy (turingtest):

that's what rhaukus did yesterday, but he got the wrong answer... ideas imperialist?

OpenStudy (akshay_budhkar):

dy=mdx + xdm

OpenStudy (anonymous):

I'm trying a couple of things, I'll let you know if anything works.

OpenStudy (unklerhaukus):

thanks

OpenStudy (akshay_budhkar):

m + x dm/dx = 4m^2 + 5 m +1

OpenStudy (akshay_budhkar):

x dm /dx= 4m^2+4m+1

OpenStudy (akshay_budhkar):

dx/x=dm/4m^2+4m+1

OpenStudy (akshay_budhkar):

um then integrate :P

OpenStudy (akshay_budhkar):

u can do from here?

OpenStudy (unklerhaukus):

ill give it a go ,

OpenStudy (anonymous):

Akshay is right, that's how you do it

OpenStudy (anonymous):

is y = -1/2 x a solution?

OpenStudy (anonymous):

a particular* solution.

OpenStudy (unklerhaukus):

\[ln(x)={1 \over 4x+2}+c\]

OpenStudy (anonymous):

\[\ln(x)=\frac{-1}{4v+2}+C\] before you get too far without the negative sign and with x instead of v

OpenStudy (anonymous):

yeah, \[y=-\frac{1}{2}x\]is one solution. just let y = mx like akshay said, and when you plug that in you get an equation for m.

OpenStudy (anonymous):

Sorry, m instead of v in what I said above, I always use v's for these types of questions!

OpenStudy (akshay_budhkar):

lol i prefer m :P

OpenStudy (unklerhaukus):

i meant \[lnx=-1/(4m+2) +c\]

OpenStudy (anonymous):

Me too! I prefer 'V'

OpenStudy (watchmath):

\(y'/x=4u^2/x+5u/x+1\) where \(u=y/x\) \(y'/x-u/x=4u^2/x+4u/x+1/x\) \(\frac{xy'-y}{x^2}=4u^2/x+4u/x+1/x\) \(du/dx=4u^2/x+4u/x+1/x\) \(du/(4u^2+4u+1)=dx/x\) \(du/(2u+1)^2=dx/x\) \(-\frac{1}{2}(2u+1)^{-1}=\ln|x|+C\) \(-\frac{1}{2}(\frac{2y}{x}+1)^{-1}=\ln|x|+C\) Let see ...

OpenStudy (akshay_budhkar):

just re-substitute the m now to get your equation.. i think we are done, aren't we?

OpenStudy (unklerhaukus):

\[{-1 \over 2}(2{y \over x}+1)^{-1}=ln|x| +C\] ... \[x+2(2y+x)ln|kx|=0\] where \[(e^C=k)\]

OpenStudy (akshay_budhkar):

i think that your answer is correct :D

OpenStudy (turingtest):

I thought you said it had an arctan in it Unk? what did you do wrong last time?

OpenStudy (akshay_budhkar):

did u try wolfphram @turing?

OpenStudy (turingtest):

I'm going by unk's last post, no haven't looked at wolf looking at this http://openstudy.com/users/unklerhaukus#/updates/4f14e000e4b0b9109c92149b

OpenStudy (watchmath):

\(-1/(\ln|x|+C) =4(y/x)+2\) \(-x/(\ln |x|+C)=4y+2x\) \(y=1/4( -x/(\ln|x|+c) -2x)\)

OpenStudy (akshay_budhkar):

@turing just check with wolfphram.. it is possible to have two different ways of represnting the same function

OpenStudy (turingtest):

I just assumed they were too different, but I suppose you are right.

OpenStudy (akshay_budhkar):

not a Guru at using wolhphram but i assume this is right http://www.wolframalpha.com/input/?i=dy%2Fdx%3D%284y%5E2%2Fx%5E2%29+%2B+5y%2Fx+%2B+1